OpenStudy (anonymous):

A spherical snowball is melting in such a way that its diameter is decreasing at rate of 0.3 cm/min. At what rate is the volume of the snowball decreasing when the diameter is 12 cm.

OpenStudy (anonymous):

Do you know an equation that represents the volume of a sphere and the diameter?

OpenStudy (anonymous):

v = 4π/3 r3 my answer is 135.71, and so is my friend's i am confused. went online to see, same answer.

OpenStudy (anonymous):

Well, the radius is half the diameter, and you want to express it in diameter, so it would be\[V = \frac{4}{3}\pi(\frac{d}{2})^3\]. Then you need to differentiate with respect to time

OpenStudy (anonymous):

first off your answer should be negative

OpenStudy (anonymous):

(Note the answer is a positive number).

OpenStudy (anonymous):

really? how is it positive if the volume is decreasing?

OpenStudy (anonymous):

A spherical snowball is melting in such a way that its diameter is decreasing at rate of 0.3 cm/min. At what rate is the volume of the snowball decreasing when the diameter is 12 cm. (Note the answer is a positive number).

OpenStudy (anonymous):

Semantics "At what rate is the volume of the snowball decreasing" I guess they mean the magnitude since they already ask for the decreasing value

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i will guess that if the diameter is decreasing at a rate of .3 cm/minute then the radius is decreasing at a rate of 0.15 cm / minute differentiate \[V=\frac{4}{3}\pi r^3\] get \[V'=4\pi r^2 r'\] replace \(r'\) by \(0.15\) and \(r\) by \(6\) and see what you get